OpenStudy (steve816):

Quick trigonometry question. Are you allowed to cross multiply to prove identities?

OpenStudy (amistre64):

generally you have to work one side to the other ... altering both sides at the same time is not a good approach.

OpenStudy (ikram002p):

Your allowed with caution for values that make denominator =0

OpenStudy (amistre64):

cross multiplication, or working backwards is a good way to discover an approach to reach the solution tho. but your proofing has to rely on the manipulation of one side into the other.

OpenStudy (steve816):

For example:\[\frac{ \cos \theta }{ \cos \theta - \sin \theta } = \frac{ 1 }{ 1 - \tan \theta }\]

OpenStudy (steve816):

Cross multiplying to prove this seems much easier.

OpenStudy (amistre64):

just divide top and bottom by cos(t) ... right?

OpenStudy (amistre64):

since tan(t) is not defined for cos(t)=0 anyways, then there is not discrepancy

OpenStudy (ikram002p):

Yeah showing cos t-sin t /cos = tan t -1 is easier

OpenStudy (steve816):

oh wow, you're right I guess.

OpenStudy (steve816):

Alright, thanks guys.